OpenStudy (aaronandyson):

If tan A + cot A = 5 find the value of tan^2 A + cot^2 A

OpenStudy (aaronandyson):

@satellite73

OpenStudy (aaronandyson):

@ParthKohli

OpenStudy (aaronandyson):

@iambatman

OpenStudy (anonymous):

\[\tan(A) + \cot(A) = 5 \\ \tan(A) + \frac{1}{\tan(A)} = 5 \\ \tan^2(A) + 1 = 5 \tan(A) \\ \tan^2(A) - 5 \tan(A) +1 = 0\]

OpenStudy (anonymous):

Solve this quadratic to find \(\tan(A)\).

OpenStudy (aaronandyson):

I don't know how to.

OpenStudy (aaronandyson):

@eta

OpenStudy (aaronandyson):

@sleepyhead314 @sleepyjess

OpenStudy (aaronandyson):

@ParthKohli

Parth (parthkohli):

Just square both sides.

OpenStudy (anonymous):

I am dumb..

OpenStudy (aaronandyson):

Can you explain it in a better way?

Parth (parthkohli):

\[\tan A + \cot A = 5\]Square both sides.

OpenStudy (anonymous):

\[(ta(A) + \cot(A))^2 = 5^2\] now proceed..

OpenStudy (aaronandyson):

25?

OpenStudy (anonymous):

*tan(A)..

OpenStudy (aaronandyson):

@ParthKohli Help me in more questions?